OpenStudy (anonymous):

find the 21st term of this sequence: -4, -1.5, 1, 3.5,....

OpenStudy (anonymous):

From the given, Let a and b be the missing variable of the equation, so: y=ax+b, from the given (1) 3a+b=14 (2) 8a+b=-3.5 (2) - (1) should give you 5a=-17.5, so that a=-3.5 Substituting a=-3.5 in (1), 3(-3.5)+b=14 which gives b=24.5 So the equation is y=-3.5x+24.5

OpenStudy (anonymous):

may i ask why u turned 3.5 negative? @trumpetmaster7777

ganeshie8 (ganeshie8):

looks @trumpetmaster7777 posted reply for a different question in this post hmm

OpenStudy (anonymous):

yah cause i dont get it..lol

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

are we humming?

ganeshie8 (ganeshie8):

lol ok lets do some math now

ganeshie8 (ganeshie8):

you're doing sequence problems, right @ciara_xo

OpenStudy (anonymous):

yes sir :p

ganeshie8 (ganeshie8):

-4, -1.5, 1, 3.5,....

ganeshie8 (ganeshie8):

first term a = ?

OpenStudy (anonymous):

-4?

ganeshie8 (ganeshie8):

right

ganeshie8 (ganeshie8):

common difference d = 3.5-1 = ?

OpenStudy (anonymous):

3.4?

ganeshie8 (ganeshie8):

how the heck..

ganeshie8 (ganeshie8):

3.5 - 1 = 2.5

ganeshie8 (ganeshie8):

come back to math ciara

OpenStudy (anonymous):

sorry im half asleep rn :/

ganeshie8 (ganeshie8):

i can see that :/

OpenStudy (anonymous):

i had a rough night too lol but ill snap out of it.

ganeshie8 (ganeshie8):

ok, so we have a = -4 d = 2.5

ganeshie8 (ganeshie8):

\(\large a_n = a+(n-1)d\)

ganeshie8 (ganeshie8):

you wanto find 21st term, so put n = 21

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

as you might be familiar with that above, that gives the nth term in arithmetic sequence

ganeshie8 (ganeshie8):

\(a_{21} = -4 + (21-1)*2.5\) = ?

OpenStudy (anonymous):

after the parenthesis is that a plus sign?

OpenStudy (anonymous):

or multiply?

ganeshie8 (ganeshie8):

* = multiply + = plus

OpenStudy (anonymous):

69=fun time

OpenStudy (anonymous):

@trumpetmaster7777 can i help you or your just being immature?

OpenStudy (anonymous):

can't help

OpenStudy (anonymous):

just a dog

ganeshie8 (ganeshie8):

what do you get ciara as final answer

OpenStudy (anonymous):

and @ganeshie8 its cause i cant really see it on my computer but im trying to solve it right now

ganeshie8 (ganeshie8):

ok its a multiply sign.

OpenStudy (anonymous):

ok lemme try it out :)

OpenStudy (anonymous):

i got 54..i hope thats right lol

ganeshie8 (ganeshie8):

check again

ganeshie8 (ganeshie8):

it seems you added 4 to 50, instead of subtracting !

OpenStudy (anonymous):

yes thats what i did -.- so its 46?

ganeshie8 (ganeshie8):

\(\color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you for your help !

ganeshie8 (ganeshie8):

np, yw !